OpenStudy (anonymous):

Would 3sqrt(3x^2) simplify to 3sqrt(3)? I'm just trying to clarify that since x^0=1 it would simply cancel out. Thanks!

OpenStudy (anonymous):

\[3\sqrt{3x ^{2}}=3\sqrt{3}x\]

OpenStudy (anonymous):

would you also be able to write it 3x\[\sqrt{3}\] ?

OpenStudy (anonymous):

\[3x \sqrt{3}\]